A​ 9-year-old girl did a science fair experiment in which she tested professional touch therapists to see if they could sense her energy field. She flipped a coin to select either her right hand or her left​ hand, and then she asked the therapists to identify the selected hand by placing their hand just under her hand without seeing it and without touching it. Among 258 ​trials, the touch therapists were correct 113 times. Use a 0.10 significance level to test the claim that touch therapists use a method equivalent to random guesses. Do the results suggest that touch therapists are​ effective?

a. IDENTIFY THE NULL AND ALTERNATIVE HYPOTHESES FOR THIS TEST

B. TEST STATISTIC=

C. P VALUE =

D. IDENTIFY THE CONCLUSION FOR THIS HYPOTHESIS TEST
Answer #1

X = 113 , n = 258 , p = X/n = 0.438 , alpha = 0.10

A) Ho : P = 0.5 vs. Ha : P not equal 0.5

This is Z test one population proportion test

two tailed z test

B) test statistics

Zstat = ( p - P ) / √ [ ( P * ( 1 - P ) ) ÷ n ] = -1.992

|Z stat |=| 1.992 | > 1.645 = Zcritical

We reject Ho

C) P value :- 2 * P ( Zc > |1.992| ) = 0.463

P value = 0.0463 < 0.10 = alpha

We reject null hypothesis

D) conclusion : there is sufficient evidence to support claim that the touch therapists are effective at alpha 0.10